Vin C Vout Vm= V2 2 45°(V), ZR=1.2, 2c= 1/; N. Determine iz and Vout (engineering notation, Z20). Notes Comments 2 pa
Answer #1

- Poul Griven , vin = r Luso ZR = in To find IR = win = in ZR+Z ve Lugo - ; 12 Lus inj = V Lugº (j+1) = 6.71 +0.713)cj+) = 0.че | lay och+ss cго «» Ne = It - 1 3 o°. = -1 Lss = -(сеѕ о т т S1 • 4 ) 2 - ( ) =